### v1

# Assessing the Robustness of Retrieval-Augmented Generation Systems in K-12 Educational Question Answering with Knowledge Discrepancies

Abstract
Retrieval-Augmented Generation (RAG) systems have demonstrated remarkable potential as question answering systems in the K-12 Education domain, where knowledge is typically queried within the restricted scope of authoritative textbooks. However, the discrepancy between textbooks and the parametric knowledge in Large Language Models (LLMs) could undermine the effectiveness of RAG systems. To systematically investigate the robustness of RAG systems under such knowledge discrepancies, we present EduKDQA, a question answering dataset that simulates knowledge discrepancies in real applications by applying hypothetical knowledge updates in answers and source documents. EduKDQA includes 3,005 questions covering five subjects, under a comprehensive question typology from the perspective of context utilization and knowledge integration. We conducted extensive experiments on retrieval and question answering performance. We find that most RAG systems suffer from a substantial performance drop in question answering with knowledge discrepancies, while questions that require integration of contextual knowledge and parametric knowledge pose a challenge to LLMs. All resources will be released to foster further research.
[…]
Assessing the Robustness of Retrieval-Augmented Generation Systems in K-12 Educational Question Answering with Knowledge Discrepancies
Tianshi Zheng††thanks: Equal Contribution♠, Weihan Li11footnotemark: 1♣, Jiaxin Bai♠, Weiqi Wang♠, Yangqiu Song♠ ♠Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China ♣Department of Mechanical Engineering, The University of Tokyo, Tokyo, Japan tzhengad@connect.ust.hk
| Question Type | Reasoning Pattern and Example Question | Hypothetical Knowledge Update | ||
| Simple Direct |
| What type of light is detected by night vision goggles? |
| Multi-hop Direct |
| Which scientist developed an equation that can calculate the wavelength of a particle? |
| Multi-hop Distant |
| Which pump creates an electrochemical gradient that enables secondary active transport to occur? |
| Multi-hop Implicit |
| Who discovered the radioactive element that is commonly found in uranium ores? |
| Distant Implicit |
| Who discovered the organelle that is responsible for the biological process that produces ATP? |
1 Introduction
In K-12 education, Question Answering (QA) systems serve as an important resource for learning assistance, where answers are precisely provided within a restricted knowledge scope from authoritative sources (i.e., textbooks) Raamadhurai et al. (2019); Soares et al. (2021). Meanwhile, benefiting from the emergent abilities Wei et al. (2022) of Large Language Models (LLMs) and advanced information retrieval (IR) methods, Retrieval-Augmented Generation (RAG) systems have achieved remarkable performance in various knowledge-intensive tasks in natural language processing Lewis et al. (2020); Jiang et al. (2023); Gao et al. (2024), demonstrating their great potential as QA systems in K-12 education Gan et al. (2023); Kasneci et al. (2023); Yan et al. (2024).
In K-12 educational QA, one of the primary concerns is ensuring that the knowledge conveyed in the answer is consistent with the officially designated textbooks Extance (2023). However, there are notable discrepancies between the knowledge in textbooks and the internal knowledge of LLMs, due to the evolving nature of facts Arbesman (2012), updates in pedagogical approaches Provenzo et al. (2011), as well as regional and cultural variations in content Patel (2015). It remains unclear whether RAG systems can robustly incorporate knowledge from authoritative sources and generate consistent answers under such knowledge discrepancies (scenario illustrated in Figure 1).
To fill this gap, we aim to systematically assess the robustness of RAG systems in performing question answering in K-12 education when encountering knowledge discrepancies. We present EduKDQA (Educational Knowledge Discrepancy Question Answering), a new dataset containing 3,005 multiple-choice questions covering the subjects of Physics, Chemistry, Biology, Geography, and History from the middle-school curriculum. To simulate the knowledge discrepancy between LLMs and textbooks, we conduct a hypothetical knowledge update, in which we modify the original factual knowledge from the textbooks into plausible alternatives while maintaining coherent and consistent context. Moreover, we tailored a comprehensive question typology to stress-test the context utilization and knowledge integration abilities of LLMs under such scenarios.
We conducted extensive experiments with various retrieval methods and LLMs. We find that most RAG systems still suffer from a considerable performance drop when facing knowledge discrepancies. Notably, while most LLMs can incorporate distant contextual facts well, they struggle to integrate their parametric knowledge with contextual knowledge effectively. In terms of retrieval, traditional lexical-based methods show advantages due to specificity of the academic terms, and their performance may be further enhanced through an ensemble reranking mechanism. To encourage further research, we will make our benchmark dataset and the associated code publicly available.
2 The EduKDQA Dataset
In this section, we introduce the methodology of hypothetical knowledge update and the design of our question typology. The curation pipeline and detailed statistics of the EduKDQA dataset are provided in Appendix B and E, respectively.
2.1 Hypothetical Knowledge Update
One of the core objectives of our dataset is to simulate the knowledge discrepancy between LLMs and authoritative textbooks when performing educational QA. However, such discrepancies are often fuzzy and highly sparse in real-world data, making it infeasible to collect and organize. Consequently, we designed the methodology of hypothetical knowledge update, performing it on high-quality open-source textbooks. The general procedures are as follows: 1) Curate factual questions from textbook paragraphs following our designed question typology. 2) Select a plausible but factually incorrect answer as the updated ground-truth answer. 3) Replace all occurrences of the original answer in the paragraph with the updated answer and adjusted other relevant statements in the context to ensure that the updated paragraph is coherent and consistent. This process is further guaranteed through extensive human curation and verification. Examples of hypothetical knowledge updates in our dataset are provided in Appendix D.
2.2 Question Typology
We identify two potential challenges for LLMs when performing QA with knowledge discrepancy: 1) Context Utilization: Whether LLMs can identify and utilize the corresponding facts from the context; and 2) Knowledge Integration: Whether LLMs can incorporate their own parametric knowledge with the contextual ones in question answering. To this end, we designed our question typology, as illustrated in Table 1, to investigate such abilities. Based on the two basic question types, Simple Direct and Multi-hop Direct, we developed the Multi-hop Distant type to evaluate the context utilization ability for distant facts from the passage, and the Multi-hop Implicit111The word ’implicit’ indicates that the questions indirectly query the updated facts by embedding them within the middle of the multi-hop reasoning chain. type to evaluate the knowledge integration ability that combines their own factual knowledge with retrieved ones. Moreover, the Distant Implicit type poses a greater challenge by combining both features. To ensure our evaluation of knowledge integration ability is independent of knowledge coverage, we restrict the facts requiring LLMs’ own knowledge to be head-knowledge only Sun et al. (2024), which we believe are well covered by all tested LLMs.
3 Experiments and Analyses
Typically, RAG systems first conduct document retrieval based on given queries, then perform question answering with LLMs based on the retrieved information loaded in the context. In this section, we comprehensively evaluate and analyze the performance of retrieval methods and LLMs on the EduKDQA dataset. For details of all tested methods and models, please refer to Appendix C.
| Retrieval Methods | Category | R@1 | R@5 |
| TF-IDF Spärck Jones (1972) | Lexical | 65.82 | 88.72 |
| BM25 Robertson et al. (1994) | Lexical | 82.73 | 95.27 |
| SPLADE Formal et al. (2021) | Lexical/Dense | 78.04 | 90.12 |
| Contriever Izacard et al. (2022) | Dense | 53.18 | 81.80 |
| Con.-msmarco Izacard et al. (2022) | Dense | 76.17 | 93.54 |
| Mistral-embed Mistral (2023a) | Dense | 78.74 | 95.31 |
| Ada-002 OpenAI (2022) | Dense | 79.23 | 95.44 |
| Query Rewrite Ma et al. (2023) | Pre-Retrieval | 78.87 | 94.21 |
[…]
| Con.-msmarco (fine-tuned) | Dense+FT | 87.95 | 99.50 |
3.1 Retrieval Performance
Experimental results of the retrieval methods are presented in Table 2. Traditional lexical retrieval methods, such as BM25, demonstrated strong performance on our dataset, while dense retrieval methods, such as Mistral-embed and Ada-002, achieved comparable performance. Since our dataset focuses on the K-12 education domain, lexical retrieval effectively captures domain-specific keywords in the queries and identifies the corresponding documents. This characteristic underscores the need for fine-tuning dense retrieval models on our dataset. To this end, we fine-tuned both Contriever and Contriever-msmarco on our documents, resulting in significant improvements and highlighting the importance of corpus-specific fine-tuning in educational document retrieval.
Moreover, ensemble methods have demonstrated their effectiveness in various information retrieval tasks Thakur et al. (2021). We implemented a hybrid approach that retrieves the top k222We use k = 6 as the optimal hyperparameter setting. documents with Ada-002, followed by re-ranking with BM25. This ensemble method marginally improved the retrieval performance in both metrics. Conversely, query rewriting333We use Mistral-small-2409 as the LLM query rewriter. Ma et al. (2023) implemented upon BM25 did not yield performance gains. Detailed results across subjects and question types are provided in Appendix F.
| Large Language Models | Question Typology | Average | ||||
| Simple Direct | Multi-hop Direct | Multi-hop Distant | Multi-hop Implicit | Distant Implicit | ||
| Mistral-7b Mistral (2023b) | 77.70 | 69.31 | 72.74 | 45.32 | 33.98 | 61.26 |
| Mixtral-8x22b Mistral (2023c) | 84.10 | 84.58 | 87.15 | 73.86 | 65.37 | 79.67 |
| Mistral-small-2409 Mistral (2024a) | 87.87 | 88.70 | 89.69 | 72.18 | 60.81 | 80.77 |
| Mistral-large-2407 Mistral (2024b) | 83.93 | 83.51 | 87.29 | 82.25 | 70.57 | 81.66 |
| Gemini-1.5-flash Google (2024a) | 80.98 | 82.44 | 87.57 | 76.02 | 63.25 | 78.54 |
| Gemini-1.5-pro Google (2024b) | 86.56 | 87.63 | 88.42 | 76.26 | 63.58 | 81.10 |
| Llama3-8b Meta (2024) | 90.33 | 88.85 | 89.69 | 63.55 | 49.43 | 77.77 |
[…]
| GPT-3.5-turbo OpenAI (2022) | 92.62 | 90.53 | 91.24 | 71.22 | 57.72 | 81.73 |
| GPT-4 OpenAI (2023a) | 89.51 | 89.47 | 90.68 | 78.66 | 70.89 | 84.46 |
| GPT-4-turbo OpenAI (2023b) | 95.74 | 96.18 | 96.19 | 81.06 | 71.71 | 88.99 |
| GPT-4o OpenAI (2024a) | 91.97 | 94.81 | 93.64 | 81.29 | 70.41 | 87.09 |
| Claude-3-sonnet Anthropic (2024a) | 94.59 | 92.82 | 93.64 | 77.70 | 60.65 | 84.69 |
[…]
| o1-preview OpenAI (2024b) | 95.08 | 97.71 | 97.46 | 86.33 | 78.86 | 91.68 |
3.2 Question Answering Performance
Experimental results of LLMs in question answering are presented in Table 3. Most LLMs exhibited comparable performance for question types Simple Direct, Multi-hop Direct, and Multi-hop Distant. This suggests that both multi-hop reasoning and distant context utilization do not pose significant challenges for modern LLMs. However, for Multi-hop Implicit questions requiring the integration of contextual and internal knowledge, a substantial performance disparity emerges between smaller open-source LLMs and advanced models. This disparity is further amplified for Distant Implicit questions. The most capable model, o1-preview, attains over 80% accuracy on Implicit questions, whereas Mistral-7b’s performance falls below 40%. These findings indicate that knowledge integration is an emergent capability presenting greater difficulties for LLMs under knowledge discrepancies, particularly when coupled with the need for distant context utilization.
| RAG System | Hypothetical Knowledge Update | Drop | |
| Before | After | ||
| Llama3-8b + Ada-002 | 87.49 | 62.60 | 24.89 |
| Llama3-8b + Rerank | 88.49 | 66.02 | 22.47 |
| GPT-4o + Ada-002 | 96.57 | 69.65 | 26.92 |
| GPT-4o + Rerank | 97.10 | 73.71 | 23.39 |
3.3 Overall Performance
How do knowledge discrepancies affect the performance of RAG systems in educational application? To answer this practical question, we selected two representative LLMs: Llama3-8b from open-source models and GPT-4o from commercial models, and combined them with two high-performing retrieval methods: Ada-002 and hybrid rerank. The resulting RAG systems were tested on questions before and after hypothetical knowledge updates, with results presented in Table 4. We observed an accuracy drop of 22-27%, indicating a substantial performance degradation in modern RAG systems when faced with knowledge discrepancies.
4 Related Work
Retrieval-Augmented Generation
Educational Question Answering
Prior to the emergence of RAG systems utilizing LLMs, various educational QA systems were developed to provide pedagogically appropriate responses to student inquiries Abdi et al. (2018); Agarwal et al. (2019). While recent literature explores LLM applications in QA and learning assistance roles Nye et al. (2023); Kuo et al. (2023); Wang et al. (2024).
Knowledge Discrepancy in LLMs
Mitigating knowledge discrepancies or conflicts in applications is a fundamental challenge in LLM research Xu et al. (2024). Researchers have proposed tuning-based Li et al. (2022) and prompting-based Zhou et al. (2023) methods to enhance LLMs’ robustness under such conflicts. However, these conflicts in educational applications and RAG systems remain relatively underexplored.
5 Conclusion
This paper systematically evaluates the robustness of retrieval-augmented generation systems in K-12 educational question answering under knowledge discrepancies using a comprehensive dataset. Experimental findings reveal substantial performance degradation in RAG systems when faced with knowledge discrepancies, which is primarily attributed to deficiencies in incorporating contextual and parametric knowledge in question answering—an emergent and challenging ability for contemporary large language models.
Limitations
We discuss three main limitations of our work.
First, EduKDQA employs the approach of hypothetical knowledge updates to effectively simulate real-world knowledge discrepancies for two primary reasons: (1) Real-world knowledge conflicts are often sparse, noisy, and difficult to systematically collect or organize into a cohesive dataset suitable for large-scale evaluation. Hypothetical updates provide a scalable and controlled alternative, allowing us to bypass these limitations. (2) By leveraging a systematic annotation and curation pipeline, we can generate questions with diverse and well-defined reasoning patterns that align with our typology, enabling more robust evaluation of complex question-answering tasks. For future research centered on real-world knowledge discrepancies, we recommend an alternative methodology that incorporates temporal attributes Chen et al. (2021); Zhang and Choi (2024). This approach focuses on identifying outdated facts and capturing time-sensitive data (e.g., economic trends, annual events, or societal changes) to construct datasets that reflect real-world knowledge updates. While promising, this method is constrained by the limited overlap between time-sensitive data—often numerical or attribute-specific—and the broader contextual needs of educational question-answering tasks, which may reduce the comprehensiveness of the resulting datasets.
Next, regarding document retrieval, some recent hierarchical retrieval paradigms, such as GraphRAG Edge et al. (2024) and HippoRAG Gutiérrez et al. (2024), are not included in our experiments due to their implementation complexity. However, we believe that such structured paradigms could effectively enhance retrieval performance in our scenario, as educational documents are well-structured and contain high-quality factual knowledge.
Finally, this paper primarily evaluates the robustness of RAG systems in the proposed scenario, with experiments conducted using various retrieval methods and large language models. Potential improvements could be achieved through the design of tailored reasoning frameworks via prompting, in-context learning or alignment in LLMs, which we leave for future research.
Ethics Statement
In constructing the EduKDQA dataset, we collected text from open-access textbooks. Detailed sources and licenses are provided in Appendix A. The human curation and verification in our annotation pipeline were carried out by a group of postgraduate students with extensive experience in NLP research. We ensured that the updated knowledge is free from harmful or toxic content. It is important to note that our dataset is designed solely to evaluate the robustness of Retrieval-Augmented Generation systems under scenarios with knowledge discrepancies and is not suitable for assessing the factual accuracy of QA systems.
Appendix A Source Textbooks
The documents in the EduKDQA dataset is organized based on the following public textbooks:
Physics
Physics by Openstax (CC-BY-4.0444https://creativecommons.org/licenses/by/4.0/deed.en) https://openstax.org/details/books/physics
Chemistry
Chemistry by Openstax (CC-BY-4.011footnotemark: 1) https://openstax.org/details/books/chemistry-2e
Biology
Biology by Openstax (CC-BY-4.011footnotemark: 1) https://openstax.org/details/books/biology-2e
History
World History by OER Commons (CC-BY-NC-4.0555https://creativecommons.org/licenses/by-nc/4.0/deed.en) https://oercommons.org/courses/world-history-2
Geography
World Regional Geography by Sailor Academy (CC-BY-3.0666https://creativecommons.org/licenses/by/3.0/)
https://learn.saylor.org/course/view.php?id=722
Meanwhile, the detailed topics included in each subjects are illustrated in Figure 2.
Appendix B Curation Pipeline
The curation pipeline of our dataset is illustrated in Figure 3. We first perform triplet extraction on the textbook documents and generate a document-level knowledge graph (KG). Next, we perform sub-graph matching based on fixed reasoning patterns to sample candidate queries, and selectively transform them into natural language questions. Then, hypothetical knowledge update is executed and verified to guarantee consistency between the updated answer and the document. Context-focused question types, including Simple Direct, Multi-hop Direct, and Multi-hop Distant, are acquired through this process. To generate the other two Implicit question types that require knowledge integration, we perform extra QA augmentation followed by an expert verification process. Our data curation process is performed by a mixture of human annotators and LLMs. For LLM annotation, we adopted Claude-3.5-Sonnet for its outstanding instruction-following ability. The total API cost for data annotation is approximately 300 USD.
Appendix C Model Details
In this section, we briefly introduce all tested retrieval methods and large language models in our benchmark experiment.
In the retrieval stage of our experiments, we employ a diverse range of retrieval methods. For traditional lexical retrieval, we include TF-IDF Spärck Jones (1972), which vectorizes documents and queries based on term frequency and inverse document frequency, and BM25 Robertson et al. (1994), which enhances TF-IDF with document length normalization and probabilistic term weighting. We also incorporate SPLADE Formal et al. (2021), a method that bridges lexical and dense retrieval paradigms. For dense retrieval, we evaluate several methods that encode questions and documents into the same vector space: Contriever Izacard et al. (2022), an unsupervised text encoder; its fine-tuned variant Contriever-msmarco; and two closed-source embedding models, Mistral-embed Mistral (2023a) and Ada-002 OpenAI (2022). Additionally, we explore a pre-retrieval method, Query Rewrite Ma et al. (2023), which reformulates queries to improve retrieval performance. Finally, we implement a Hybrid Rerank approach that combines BM25 and Ada-002, leveraging the strengths of both lexical and dense retrieval methods.
In our evaluation, we employ a diverse set of state-of-the-art large language models to assess their performance across various tasks. The models include: Mistral AI’s open-source models, ranging from the compact Mistral-7b Mistral (2023b) to the more advanced MoE model Mixtral-8x22b Mistral (2023c), and their latest iterations Mistral-small-2409 Mistral (2024a) and Mistral-large-2407 Mistral (2024b). Google’s Gemini models are represented by Gemini-1.5-flash Google (2024a) and Gemini-1.5-pro Google (2024b). Meta’s Llama3 series is included with both 8b and 70b parameter versions Meta (2024). We also evaluate OpenAI’s models, including GPT-3.5-turbo OpenAI (2022), GPT-4 OpenAI (2023a), GPT-4-turbo OpenAI (2023b), and GPT-4o OpenAI (2024a). Anthropic’s LLMs are represented by Claude-3-sonnet Anthropic (2024a) and Claude-3.5-sonnet Anthropic (2024b). Lastly, we include OpenAI’s latest o1 series, o1-mini and o1-preview OpenAI (2024b), which achieve remarkable performance across various metrics through inference-time scaling. This comprehensive selection allows us to compare a wide range of model architectures and sizes, providing valuable insights into the current state of LLM capabilities.
Appendix D Dataset Example
We provide an example question and its corresponding paragraph before and after hypothetical knowledge update in Table 5.
| Original Paragraph | … The halophiles, which means "salt-loving", live in environments with high levels of salt. They have been identified in the Great Salt Lake in Utah and in the Dead Sea between Israel and Jordan, which have salt concentrations several times that of the oceans. … |
| Updated Paragraph | … The halophiles, which means "pressure-loving", live in environments with high levels of pressure. They have been identified in the Mariana Trench in western Pacific Ocean, which have higher pressure than other environments. … |
| Question | In which type of environments do halophiles typically live? A. High Acidity Environments. B. High Salt Environments. C. High Pressure Environments. D. High Sugar Environments. |
| Original Answer | B. High Salt Environments. |
| Updated Answer | C. High Pressure Environments. |
Appendix E Dataset Statistics
The statistics for the documents and questions in the EduKDQA dataset are provided in Table 6 and Table 7, respectively.
| Document Statistics | Subjects | Total | ||||
| Chem. | Bio. | Phys. | Geo. | Hist. | ||
| Average Document Length | 315.5 | 588.1 | 647.3 | 429.8 | 277.1 | 437.3 |
| Num of Documents | 291 | 671 | 166 | 471 | 606 | 2205 |
| # Questions (avg. length) | Question Types | Total | ||||
| Simple Direct | Multi-hop Direct | Multi-hop Distant | Multi-hop Implicit | Distant Implicit | ||
| Chemistry | 73 (12.4) | 88 (17.8) | 78 (17.9) | 40 (17.4) | 68 (21.1) | 347 (17.3) |
| Biology | 148 (11.1) | 170 (16.7) | 248 (16.7) | 94 (16.7) | 213 (20.7) | 873 (16.7) |
| Physics | 46 (12.0) | 49 (16.7) | 49 (15.6) | 41 (16.9) | 39 (19.9) | 224 (16.1) |
| Geography | 141 (12.1) | 147 (16.7) | 162 (16.8) | 96 (17.7) | 144 (21.3) | 690 (16.9) |
| History | 202 (12.8) | 201 (16.1) | 171 (17.0) | 146 (18.1) | 151 (21.6) | 871 (16.8) |
| Total | 610 (12.1) | 655 (16.7) | 708 (16.9) | 417 (17.5) | 615 (21.0) | 3005 (16.8) |
| Retrieval Methods | Subjects | Question Types | Average | ||||||||
| chem. | bio. | phys. | geo. | hist. | Sim. Dir. | Mul. Dir. | Mul. Dis. | Mul. Imp. | Dis. Imp. | ||
| BM25 | 87.90 | 84.65 | 74.55 | 78.26 | 81.52 | 79.34 | 85.04 | 86.30 | 75.78 | 84.23 | 82.73 |
| Mistral-embed | 84.73 | 77.09 | 83.48 | 73.48 | 80.94 | 82.13 | 79.08 | 79.38 | 80.81 | 72.85 | 78.74 |
| Ada-002 | 84.44 | 79.50 | 82.14 | 72.75 | 81.29 | 82.30 | 80.76 | 80.08 | 80.10 | 73.01 | 79.23 |
| Large Language Models | Subjects | Question Types | Average | ||||||||
| chem. | bio. | phys. | geo. | hist. | Sim. Dir. | Mul. Dir. | Mul. Dis. | Mul. Imp. | Dis. Imp. | ||
| Llama3-8b | 77.81 | 75.60 | 75.45 | 80.14 | 78.65 | 90.33 | 88.85 | 89.69 | 63.55 | 49.43 | 77.77 |
| GPT-3.5-turbo | 78.93 | 80.53 | 82.59 | 83.04 | 82.43 | 92.62 | 90.53 | 91.24 | 71.22 | 57.72 | 81.73 |
| GPT-4-turbo | 85.30 | 89.23 | 86.16 | 90.00 | 90.13 | 95.74 | 96.18 | 96.19 | 81.06 | 71.71 | 88.99 |
| Model | Prompting Method | |
| Direct Answer | Locate-and-Answer | |
| Gemini-1.5-flash | 75.21 | 78.54 |
| GPT-3.5-turbo | 73.91 | 81.73 |
| GPT-4o | 80.43 | 87.09 |
Appendix F Result Details
The comprehensive evaluation results for retrieval approaches and large language models across various subjects and question types are presented in Table 8 and Table 9, respectively. We observed that the performance of lexical retrieval methods correlates positively with question length, while dense retrieval methods exhibit an inverse relationship. This finding suggests that there is potential for developing more sophisticated ensemble methodologies that could fully leverage the strengths of both approaches.
Appendix G Prompting Approach
In our experiments, we adopt the prompting approach of Locate-and-Answer, to facilitate active acquisition of information in the context when performing question answering. We first request LLMs to identify and locate the corresponding sentence that include the knowledge for the question from the provided document, and then reason to provide its answer. According to the experimental result in Table 3, this prompting approach can effectively improve the QA performance of LLMs compared to direct answering.