### v3

# KnowShiftQA: How Robust are RAG Systems when Textbook Knowledge Shifts in K-12 Education?

Abstract
Retrieval-Augmented Generation (RAG) systems show remarkable potential as question answering tools in the K-12 Education domain, where knowledge is typically queried within the restricted scope of authoritative textbooks. However, discrepancies between these textbooks and the parametric knowledge inherent in Large Language Models (LLMs) can undermine the effectiveness of RAG systems. To systematically investigate RAG system robustness against such knowledge discrepancies, we introduce KnowShiftQA111https://github.com/HKUST-KnowComp/KnowShiftQA. This novel question answering dataset simulates these discrepancies by applying deliberate hypothetical knowledge updates to both answers and source documents, reflecting how textbook knowledge can shift. KnowShiftQA comprises 3,005 questions across five subjects, designed with a comprehensive question typology focusing on context utilization and knowledge integration. Our extensive experiments on retrieval and question answering performance reveal that most RAG systems suffer a substantial performance drop when faced with these knowledge discrepancies. Furthermore, questions requiring the integration of contextual (textbook) knowledge with parametric (LLM) knowledge pose a significant challenge to current LLMs.
[…]
KnowShiftQA: How Robust are RAG Systems when Textbook Knowledge Shifts in K-12 Education?
Tianshi Zheng††thanks: Equal Contribution♠, Weihan Li11footnotemark: 1♣, Jiaxin Bai♠, Weiqi Wang♠, Yangqiu Song♠ ♠Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China ♣Department of Mechanical Engineering, The University of Tokyo, Tokyo, Japan tzhengad@connect.ust.hk
| Question Type | Reasoning Pattern and Example Question | Hypothetical Knowledge Update | ||
| Simple Direct |
| What type of light is detected by night vision goggles? |
| Multi-hop Direct |
| Which scientist developed an equation that can calculate the wavelength of a particle? |
| Multi-hop Distant |
| Which pump creates an electrochemical gradient that enables secondary active transport to occur? |
| Multi-hop Implicit |
| Who discovered the radioactive element that is commonly found in uranium ores? |
| Distant Implicit |
| Who discovered the organelle that is responsible for the biological process that produces ATP? |
1 Introduction
In K-12 education, Question Answering (QA) systems serve as an important resource for learning assistance, where answers are precisely provided within a restricted knowledge scope from authoritative sources (i.e., textbooks) Raamadhurai et al. (2019); Soares et al. (2021). Meanwhile, benefiting from the emergent abilities Wei et al. (2022) of Large Language Models (LLMs) and advanced information retrieval (IR) methods, Retrieval-Augmented Generation (RAG) systems have achieved remarkable performance in various knowledge-intensive tasks in natural language processing Lewis et al. (2020); Jiang et al. (2023); Gao et al. (2024), demonstrating their great potential as QA systems in K-12 education Gan et al. (2023); Kasneci et al. (2023); Yan et al. (2024).
In K-12 educational QA, one of the primary concerns is ensuring that the knowledge conveyed in the answer is consistent with the officially designated textbooks Extance (2023). However, notable knowledge shifts (i.e. factuality discrepancies) may exist between the knowledge in textbooks and the internal knowledge of LLMs, due to the evolving nature of facts Arbesman (2012), updates in pedagogical approaches Provenzo et al. (2011), as well as regional and cultural variations in content Patel (2015). It remains unclear whether RAG systems can robustly incorporate knowledge from authoritative sources and generate consistent answers under such knowledge discrepancies (scenario illustrated in Figure 1).
To fill this gap, we aim to systematically assess the robustness of RAG systems in performing question answering in K-12 education when encountering knowledge shifts. We present KnowShiftQA, a new dataset containing 3,005 multiple-choice questions covering the subjects of Physics, Chemistry, Biology, Geography, and History from the middle-school curriculum. To simulate the factuality discrepancy between LLMs and textbooks, we conduct a hypothetical knowledge update, in which we modify the original factual knowledge from the textbooks into plausible alternatives while maintaining coherent and consistent context. Moreover, we tailored a comprehensive question typology to stress-test the context utilization and knowledge integration abilities of LLMs under such scenarios.
We conducted extensive experiments with various retrieval methods and LLMs. Our findings indicate that most RAG systems exhibit a considerable performance degradation under knowledge shifts. Notably, while many LLMs can effectively incorporate distant contextual facts, they struggle to seamlessly integrate their parametric knowledge with this contextual information. In terms of retrieval, traditional lexical-based methods demonstrate advantages due to the specificity of academic terms; their performance can be further enhanced through an ensemble re-ranking mechanism. Supervised retrieval methods (e.g., Contriever), when fine-tuned directly on our dataset, yield superior performance. Overall, our results highlight the fragility of current RAG systems under knowledge shifts and the challenges LLMs face in effectively combining contextual and parametric knowledge for complex question answering.
2 The KnowShiftQA Dataset
In this section, we introduce the methodology of hypothetical knowledge update and the design of our question typology. The curation pipeline and detailed statistics of the KnowShiftQA dataset are provided in Appendix B and E, respectively.
2.1 Hypothetical Knowledge Update
One of the core objectives of our dataset is to simulate the knowledge discrepancy between LLMs and authoritative textbooks when performing educational QA. However, such discrepancies are often fuzzy and highly sparse in real-world data, making it infeasible to collect and organize. Consequently, we designed the methodology of hypothetical knowledge update, performing it on high-quality open-source textbooks. The general procedures are as follows: 1) Curate factual questions from textbook paragraphs following our designed question typology. 2) Select a plausible but factually incorrect answer as the updated ground-truth answer. 3) Replace all occurrences of the original answer in the paragraph with the updated answer and adjusted other relevant statements in the context to ensure that the updated paragraph is coherent and consistent. This process is further guaranteed through extensive human curation and verification. Examples of hypothetical knowledge updates in our dataset are provided in Appendix D.
2.2 Question Typology
We identify two potential challenges for LLMs when performing QA under knowledge shifts: 1) Context Utilization: Whether LLMs can identify and utilize the corresponding facts from the context; and 2) Knowledge Integration: Whether LLMs can incorporate their own parametric knowledge with contextual facts in question answering. To this end, we designed our question typology, as illustrated in Table 1, to investigate such abilities. The two basic question types, Simple Direct and Multi-hop Direct, aim to evaluate LLMs on simple factual recall and multi-hop reasoning. Based on these two types, we developed the Multi-hop Distant type to evaluate the context utilization ability for distant facts from the passage, and the Multi-hop Implicit222The word ’implicit’ indicates that the questions indirectly query the updated facts by embedding them within the middle of the multi-hop reasoning chain. type to evaluate the knowledge integration ability that combines their own factual knowledge with retrieved facts. Moreover, the Distant Implicit type poses a greater challenge by combining both features. To ensure our evaluation of knowledge integration ability is independent of knowledge coverage, we restrict the facts requiring LLMs’ own knowledge to be head-knowledge only Sun et al. (2024).
3 Experiments and Analyses
Typically, RAG systems first conduct document retrieval based on given queries, then perform question answering with LLMs based on the retrieved information loaded in the context. In this section, we comprehensively evaluate and analyze the performance of retrieval methods and LLMs on the KnowShiftQA dataset. For details of all tested methods and models, please refer to Appendix C.
| Retrieval Methods | Category | R@1 | R@5 |
| TF-IDF Spärck Jones (1972) | Lexical | 65.82 | 88.72 |
| BM25 Robertson et al. (1994) | Lexical | 82.73 | 95.27 |
| SPLADE Formal et al. (2021) | Lexical/Dense | 78.04 | 90.12 |
| Contriever Izacard et al. (2022) | Dense | 53.18 | 81.80 |
| Con.-msmarco Izacard et al. (2022) | Dense | 76.17 | 93.54 |
| Mistral-embed Mistral (2023a) | Dense | 78.74 | 95.31 |
| Ada-002 OpenAI (2022) | Dense | 79.23 | 95.44 |
| Query Rewrite Ma et al. (2023) | Pre-Retrieval | 78.87 | 94.21 |
[…]
| Con.-msmarco (fine-tuned) | Dense+FT | 87.95 | 99.50 |
3.1 Retrieval Performance
Experimental results of the retrieval methods are presented in Table 2. Traditional lexical retrieval methods, such as BM25, demonstrated strong performance on our dataset, while dense retrieval methods, such as Mistral-embed and Ada-002, achieved comparable performance. Since our dataset focuses on the K-12 education domain, lexical retrieval effectively captures domain-specific keywords in the queries and identifies the corresponding documents. This characteristic underscores the need for fine-tuning dense retrieval models on our dataset. To this end, we fine-tuned both Contriever and Contriever-msmarco on our documents, resulting in significant improvements and highlighting the importance of corpus-specific fine-tuning in educational document retrieval.
Moreover, ensemble methods have demonstrated their effectiveness in various information retrieval tasks Thakur et al. (2021). We implemented a hybrid approach that retrieves the top k333We use k = 6 as the optimal hyperparameter setting. documents with Ada-002, followed by re-ranking with BM25. This ensemble method marginally improved the retrieval performance in both metrics. Conversely, query rewriting444We use Mistral-small-2409 as the LLM query rewriter. Ma et al. (2023) implemented upon BM25 did not yield performance gains. Detailed results across subjects and question types are provided in Appendix F.
| Large Language Models | Question Typology | Average | ||||
| Simple Direct | Multi-hop Direct | Multi-hop Distant | Multi-hop Implicit | Distant Implicit | ||
| Mistral-7b Mistral (2023b) | 77.70 | 69.31 | 72.74 | 45.32 | 33.98 | 61.26 |
| Mixtral-8x22b Mistral (2023c) | 84.10 | 84.58 | 87.15 | 73.86 | 65.37 | 79.67 |
| Mistral-small-2409 Mistral (2024a) | 87.87 | 88.70 | 89.69 | 72.18 | 60.81 | 80.77 |
| Mistral-large-2407 Mistral (2024b) | 83.93 | 83.51 | 87.29 | 82.25 | 70.57 | 81.66 |
| Gemini-1.5-flash Google (2024a) | 80.98 | 82.44 | 87.57 | 76.02 | 63.25 | 78.54 |
| Gemini-1.5-pro Google (2024b) | 86.56 | 87.63 | 88.42 | 76.26 | 63.58 | 81.10 |
| Llama3-8b Meta (2024) | 90.33 | 88.85 | 89.69 | 63.55 | 49.43 | 77.77 |
[…]
| GPT-3.5-turbo OpenAI (2022) | 92.62 | 90.53 | 91.24 | 71.22 | 57.72 | 81.73 |
| GPT-4 OpenAI (2023a) | 89.51 | 89.47 | 90.68 | 78.66 | 70.89 | 84.46 |
| GPT-4-turbo OpenAI (2023b) | 95.74 | 96.18 | 96.19 | 81.06 | 71.71 | 88.99 |
| GPT-4o OpenAI (2024a) | 91.97 | 94.81 | 93.64 | 81.29 | 70.41 | 87.09 |
| Claude-3-sonnet Anthropic (2024a) | 94.59 | 92.82 | 93.64 | 77.70 | 60.65 | 84.69 |
[…]
| o1-preview OpenAI (2024b) | 95.08 | 97.71 | 97.46 | 86.33 | 78.86 | 91.68 |
3.2 Question Answering Performance
Experimental results of LLMs in question answering are presented in Table 3. Most LLMs exhibited comparable performance for question types Simple Direct, Multi-hop Direct, and Multi-hop Distant. This suggests that both multi-hop reasoning and distant context utilization do not pose significant challenges for modern LLMs. However, for Multi-hop Implicit questions, which requires the integration of contextual and internal knowledge, a substantial performance disparity emerges between smaller open-source LLMs and advanced models. This disparity is further amplified for Distant Implicit questions, where the reasoning chains are getting more complexed. The most capable model we tested, o1-preview, attains over 80% accuracy on Implicit questions, whereas Mistral-7b’s performance falls below 40%. These findings indicate that context-memory knowledge integration is an emergent capability presenting greater difficulties for LLMs under knowledge discrepancies, particularly when coupled with the need for distant context utilization.
| RAG System | Hypothetical Knowledge Update | Drop | |
| Before | After | ||
| Llama3-8b + Ada-002 | 87.49 | 62.60 | 24.89 |
| Llama3-8b + Rerank | 88.49 | 66.02 | 22.47 |
| GPT-4o + Ada-002 | 96.57 | 69.65 | 26.92 |
| GPT-4o + Rerank | 97.10 | 73.71 | 23.39 |
3.3 Overall Performance
How do knowledge shifts affect the performance of RAG systems in educational question-answering applications? To answer this question, we selected two representative LLMs: Llama3-8b from open-source models and GPT-4o from proprietary models. These were combined with two high-performing retrieval methods: Ada-002 and hybrid rerank. The resulting RAG systems were tested on questions before and after hypothetical knowledge updates, with the results presented in Table 4. We observed a significant accuracy drop of 22–27% after these knowledge shifts, indicating substantial performance degradation in modern RAG systems when faced with such scenarios.
4 Related Work
Retrieval-Augmented Generation
Following the categorization by Gao et al. (2024), the RAG methods employed in our experiment fall into the categories of Naive RAG (lexical, dense) and Advanced RAG (rerank, rewrite). Recently, Modular RAG has emerged to enhance the adaptability and versatility of RAG systems Shao et al. (2023); Asai et al. (2023). Furthermore, recent advances in vector databases Han et al. (2023) and neural graph databases Bai et al. (2025b) highlight the potential for adaptive and generalizable RAG search with structured data Deng et al. (2024); Bai et al. (2025a).
Educational Question Answering
Prior to the emergence of RAG systems utilizing LLMs, various educational QA systems were developed to provide pedagogically appropriate responses to student inquiries Abdi et al. (2018); Agarwal et al. (2019). While recent literature explores LLM applications in QA and learning assistance roles Nye et al. (2023); Kuo et al. (2023); Wang et al. (2024), it is also crucial to leverage logical inference abilities, such as abductive reasoning Bai et al. (2024); Yim et al. (2024); Zheng et al. (2025b) and inductive reasoning Qiu et al. (2024); Li et al. (2025); Zheng et al. (2025a), for understanding students’ learning outcomes in response to their queries.
Knowledge Discrepancy in LLMs
Mitigating knowledge discrepancies or conflicts within LLM applications is a fundamental challenge in LLM research Xu et al. (2024). Researchers have proposed tuning-based Li et al. (2022) and prompting-based Zhou et al. (2023) methods to enhance LLMs’ robustness to such conflicts. However, the implications of these conflicts in educational applications and RAG systems remain relatively underexplored. Moreover, it is important to consider the trade-off between factuality robustness Yu et al. (2024); Zong et al. (2024); Zheng et al. (2025c) and flexibility in instruction following when such discrepancies are present.
5 Conclusion
This paper systematically evaluates the robustness of RAG systems in K-12 educational question answering under knowledge discrepancies using a comprehensive dataset KnowShiftQA. Experimental findings reveal substantial performance degradation in RAG systems when faced with knowledge discrepancies, which is primarily attributed to deficiencies in incorporating contextual and parametric knowledge in question answering—an emergent and challenging ability for modern large language models.
Limitations
We discuss three main limitations of our work.
First, KnowShiftQA employs the approach of hypothetical knowledge updates to effectively simulate real-world knowledge discrepancies for two primary reasons: (1) Real-world knowledge conflicts are often sparse, noisy, and difficult to systematically collect or organize into a cohesive dataset suitable for large-scale evaluation. Hypothetical updates provide a scalable and controlled alternative, allowing us to bypass these limitations. (2) By leveraging a systematic annotation and curation pipeline, we can generate questions with diverse and well-defined reasoning patterns that align with our typology, enabling more robust evaluation of complex question-answering tasks. For future research centered on real-world knowledge discrepancies, we recommend an alternative methodology that incorporates temporal attributes Chen et al. (2021); Zhang and Choi (2024). This approach focuses on identifying outdated facts and capturing time-sensitive data (e.g., economic trends, annual events, or societal changes) to construct datasets that reflect real-world knowledge updates. While promising, this method is constrained by the limited overlap between time-sensitive data—often numerical or attribute-specific—and the broader contextual needs of educational question-answering tasks, which may reduce the comprehensiveness of the resulting datasets.
Next, regarding document retrieval, some recent hierarchical retrieval paradigms, such as GraphRAG Edge et al. (2024) and HippoRAG Gutiérrez et al. (2024), are not included in our experiments due to their implementation complexity. However, we believe that such structured paradigms could effectively enhance retrieval performance in our scenario, as educational documents are well-structured and contain high-quality factual knowledge.
Finally, this paper primarily evaluates the robustness of RAG systems in the proposed scenario, with experiments conducted using various retrieval methods and large language models. Potential improvements could be achieved through the design of tailored reasoning frameworks via prompting, in-context learning or alignment in LLMs, which we leave for future research.
Ethics Statement
In constructing the KnowShiftQA dataset, we collected text from open-access textbooks. Detailed sources and licenses are provided in Appendix A. The human curation and verification in our annotation pipeline were carried out by a group of postgraduate students with extensive experience in NLP research. We ensured that the updated knowledge is free from harmful or toxic content. It is important to note that our dataset is designed solely to evaluate the robustness of Retrieval-Augmented Generation systems under scenarios with knowledge discrepancies and is not suitable for assessing the factual accuracy of QA systems.
Acknowledgments
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong SAR, China.
Appendix A Source Textbooks
The documents in the KnowShiftQA dataset is organized based on the following public textbooks (the detailed topics included in each subjects are illustrated in Figure 2.):
Physics
Physics by Openstax (CC-BY-4.0555https://creativecommons.org/licenses/by/4.0/deed.en) https://openstax.org/details/books/physics
Chemistry
Chemistry by Openstax (CC-BY-4.011footnotemark: 1) https://openstax.org/details/books/chemistry-2e
Biology
Biology by Openstax (CC-BY-4.011footnotemark: 1) https://openstax.org/details/books/biology-2e
History
World History by OER Commons (CC-BY-NC-4.0666https://creativecommons.org/licenses/by-nc/4.0/deed.en) https://oercommons.org/courses/world-history-2
Geography
World Regional Geography by Sailor Academy (CC-BY-3.0777https://creativecommons.org/licenses/by/3.0/)
https://learn.saylor.org/course/view.php?id=722
Appendix B Curation Pipeline
The curation pipeline of our dataset is illustrated in Figure 3. We first perform triplet extraction on the textbook documents and generate a document-level knowledge graph (KG). Next, we perform sub-graph matching based on fixed reasoning patterns to sample candidate queries, and selectively transform them into natural language questions. Then, hypothetical knowledge update is executed and verified to guarantee consistency between the updated answer and the document.
Context-focused question types, including Simple Direct, Multi-hop Direct, and Multi-hop Distant, are acquired through this process. For Multi-hop Distant questions, we leverage distant facts, defined as connected triplet pairs that are separated in the document’s sequential ordering. These questions are only assigned to documents containing no fewer than 200 words. To generate the other two Implicit question types that require knowledge integration, we perform extra QA augmentation followed by an expert verification process. The design of our question typology is motivated by studies in complex logical reasoning Bai et al. (2023); Zheng et al. (2024, 2025d).
Our data curation process is performed through an integrated framework involving both human annotators and LLMs. For LLM annotation, we adopted Claude-3.5-Sonnet for its outstanding instruction-following ability. Following manual verification, 90.5% of these queries were retained or underwent minor refinements to become high-quality questions, yielding an overall success rate of 86.4%. The total API cost for data annotation is approximately 300 USD.
Appendix C Model Details
In this section, we briefly introduce all tested retrieval methods and large language models in our benchmark experiment.
In the retrieval stage of our experiments, we employ a diverse range of retrieval methods. For traditional lexical retrieval, we include TF-IDF Spärck Jones (1972), which vectorizes documents and queries based on term frequency and inverse document frequency, and BM25 Robertson et al. (1994), which enhances TF-IDF with document length normalization and probabilistic term weighting. We also incorporate SPLADE Formal et al. (2021), a method that bridges lexical and dense retrieval paradigms. For dense retrieval, we evaluate several methods that encode questions and documents into the same vector space: Contriever Izacard et al. (2022), an unsupervised text encoder; its fine-tuned variant Contriever-msmarco, which we further enhanced by applying contrastive learning Izacard et al. (2022) to fine-tune both models on the KnowShiftQA dataset for improved retrieval capability; and two closed-source embedding models, Mistral-embed Mistral (2023a) and Ada-002 OpenAI (2022). Additionally, we explore a pre-retrieval method, Query Rewrite Ma et al. (2023), which reformulates queries to improve retrieval performance. Finally, we implement a Hybrid Rerank approach that combines BM25 and Ada-002, leveraging the strengths of both lexical and dense retrieval methods.
In our evaluation, we employ a diverse set of state-of-the-art large language models to assess their performance across various tasks. The models include: Mistral AI’s open-source models, ranging from the compact Mistral-7b Mistral (2023b) to the more advanced MoE model Mixtral-8x22b Mistral (2023c), and their latest iterations Mistral-small-2409 Mistral (2024a) and Mistral-large-2407 Mistral (2024b). Google’s Gemini models are represented by Gemini-1.5-flash Google (2024a) and Gemini-1.5-pro Google (2024b). Meta’s Llama3 series is included with both 8b and 70b parameter versions Meta (2024). We also evaluate OpenAI’s models, including GPT-3.5-turbo OpenAI (2022), GPT-4 OpenAI (2023a), GPT-4-turbo OpenAI (2023b), and GPT-4o OpenAI (2024a). Anthropic’s LLMs are represented by Claude-3-sonnet Anthropic (2024a) and Claude-3.5-sonnet Anthropic (2024b). Lastly, we include OpenAI’s latest o1 series, o1-mini and o1-preview OpenAI (2024b), which achieve remarkable performance across various metrics through inference-time scaling. This comprehensive selection allows us to compare a wide range of model architectures and sizes, providing valuable insights into the current state of LLM capabilities.
Appendix D Dataset Example
We provide an example question and its corresponding paragraph before and after hypothetical knowledge update in Table 5.
| Original Paragraph | … The halophiles, which means "salt-loving", live in environments with high levels of salt. They have been identified in the Great Salt Lake in Utah and in the Dead Sea between Israel and Jordan, which have salt concentrations several times that of the oceans. … |
| Updated Paragraph | … The halophiles, which means "pressure-loving", live in environments with high levels of pressure. They have been identified in the Mariana Trench in western Pacific Ocean, which have higher pressure than other environments. … |
| Question | In which type of environments do halophiles typically live? A. High Acidity Environments. B. High Salt Environments. C. High Pressure Environments. D. High Sugar Environments. |
| Original Answer | B. High Salt Environments. |
| Updated Answer | C. High Pressure Environments. |
Appendix E Dataset Statistics
The statistics for the documents and questions in the KnowShiftQA dataset are provided in Table 6 and Table 7, respectively.
| Document Statistics | Subjects | Total | ||||
| Chem. | Bio. | Phys. | Geo. | Hist. | ||
| Average Document Length | 315.5 | 588.1 | 647.3 | 429.8 | 277.1 | 437.3 |
| Num of Documents | 291 | 671 | 166 | 471 | 606 | 2205 |
| # Questions (avg. length) | Question Types | Total | ||||
| Simple Direct | Multi-hop Direct | Multi-hop Distant | Multi-hop Implicit | Distant Implicit | ||
| Chemistry | 73 (12.4) | 88 (17.8) | 78 (17.9) | 40 (17.4) | 68 (21.1) | 347 (17.3) |
| Biology | 148 (11.1) | 170 (16.7) | 248 (16.7) | 94 (16.7) | 213 (20.7) | 873 (16.7) |
| Physics | 46 (12.0) | 49 (16.7) | 49 (15.6) | 41 (16.9) | 39 (19.9) | 224 (16.1) |
| Geography | 141 (12.1) | 147 (16.7) | 162 (16.8) | 96 (17.7) | 144 (21.3) | 690 (16.9) |
| History | 202 (12.8) | 201 (16.1) | 171 (17.0) | 146 (18.1) | 151 (21.6) | 871 (16.8) |
| Total | 610 (12.1) | 655 (16.7) | 708 (16.9) | 417 (17.5) | 615 (21.0) | 3005 (16.8) |
Appendix F Result Details
The comprehensive evaluation results for retrieval approaches and large language models across various subjects and question types are presented in Table 8 and Table 9, respectively. We observed that the performance of lexical retrieval methods correlates positively with question length, while dense retrieval methods exhibit an inverse relationship. This finding suggests that there is potential for developing more sophisticated ensemble methodologies that could fully leverage the strengths of both approaches.
| Retrieval Methods | Subjects | Question Types | Average | ||||||||
| chem. | bio. | phys. | geo. | hist. | Sim. Dir. | Mul. Dir. | Mul. Dis. | Mul. Imp. | Dis. Imp. | ||
| BM25 | 87.90 | 84.65 | 74.55 | 78.26 | 81.52 | 79.34 | 85.04 | 86.30 | 75.78 | 84.23 | 82.73 |
| Mistral-embed | 84.73 | 77.09 | 83.48 | 73.48 | 80.94 | 82.13 | 79.08 | 79.38 | 80.81 | 72.85 | 78.74 |
| Ada-002 | 84.44 | 79.50 | 82.14 | 72.75 | 81.29 | 82.30 | 80.76 | 80.08 | 80.10 | 73.01 | 79.23 |
| Large Language Models | Subjects | Question Types | Average | ||||||||
| chem. | bio. | phys. | geo. | hist. | Sim. Dir. | Mul. Dir. | Mul. Dis. | Mul. Imp. | Dis. Imp. | ||
| Llama3-8b | 77.81 | 75.60 | 75.45 | 80.14 | 78.65 | 90.33 | 88.85 | 89.69 | 63.55 | 49.43 | 77.77 |
| GPT-3.5-turbo | 78.93 | 80.53 | 82.59 | 83.04 | 82.43 | 92.62 | 90.53 | 91.24 | 71.22 | 57.72 | 81.73 |
| GPT-4-turbo | 85.30 | 89.23 | 86.16 | 90.00 | 90.13 | 95.74 | 96.18 | 96.19 | 81.06 | 71.71 | 88.99 |
Appendix G Prompting Approach
In our experiments, we adopt the prompting approach of Locate-and-Answer, to facilitate active acquisition of information in the context when performing question answering. We first request LLMs to identify and locate the corresponding sentence that include the knowledge for the question from the provided document, and then reason to provide its answer. According to the experimental result in Table 10, this prompting approach can effectively improve the QA performance of LLMs compared to direct answering.
| Model | Prompting Method | |
| Direct Answer | Locate-and-Answer | |
| Gemini-1.5-flash | 75.21 | 78.54 |
| GPT-3.5-turbo | 73.91 | 81.73 |
| GPT-4o | 80.43 | 87.09 |
Appendix H Calibration, Parametric Knowledge, and Instruction Following in RAG Systems
In our study, retrieved documents serve as the unequivocal reference, mirroring the pedagogical approach in K-12 education where textbooks are authoritative. However, our evaluation results reveal counterintuitive performance patterns in state-of-the-art LLMs, particularly concerning simple direct questions. Stronger models, such as GPT-4 and GPT-4o, occasionally underperform compared to weaker ones on these questions, which primarily assess direct factual recall from the provided (and hypothetically updated) context. This suggests that the calibration of these advanced LLMs might lead them to exhibit excessive confidence in their internal, parametric knowledge, thereby inhibiting their consistent reliance on the external documents explicitly provided as the ground truth for the task.
This observation highlights a critical tension inherent in RAG systems: the trade-off between an LLM’s adherence to its ingrained parametric knowledge (what might be termed its internal "factuality calibration") and its capacity for robust instruction following. While an LLM prioritizing its internal knowledge, especially if it aligns with generally accepted facts, could be interpreted as a form of "honesty," this perspective shifts within the specific operational context of RAG systems like those evaluated with KnowShiftQA. In such systems, LLMs are explicitly instructed to base their responses on the provided contextual documents. Therefore, when a model disregards this instruction and favors its parametric knowledge—particularly when the context contains deliberate, hypothetical updates—it signifies a limitation in its instruction-following capability rather than a commendable adherence to broader factual accuracy.
The underperformance on simple direct questions can thus be partly attributed to this conflict: the models may struggle to override a strong internal "prior" when explicitly instructed to use new, conflicting information from the context. This distinction is paramount for developing robust educational QA systems. Such systems must reliably reflect the content of designated authoritative sources (e.g., textbooks), even if these sources present information that has been updated or differs from an LLM’s training data. Consequently, the ability to flexibly integrate and prioritize instructed contextual information is a fundamental requirement for their effectiveness in dynamic knowledge environments. While in real-world scenarios with less clearly defined ground truths or more ambiguous knowledge conflicts, the sophisticated reasoning of stronger LLMs might offer advantages in navigating discrepancies, the specific demands of an educational RAG system underscore the critical importance of faithful instruction adherence.